10) how to solve x+2y=8 with the substitution of x+3y=12
Answer by Jstrasner(112)

Hey,
So for this one we need to isolate one of the variables:
x+2y=8 => x=8-2y
Then substitute 8-2y for x in the second equation:
(8-2y) + 3y = 12
then subtract 8 from both sides and you get:
y = 4
I hope this helps!